# Interaction Screening and Pseudolikelihood Approaches for Tensor Learning in Ising Models

### v1

@@ Abstract.
In this paper, we study two well known methods of Ising structure learning, namely the pseudolikelihood approach and the interaction screening approach, in the context of tensor recovery in $k$-spin Ising models. We show that both these approaches, with proper regularization, retrieve the underlying hypernetwork structure using a sample size logarithmic in the number of network nodes, and exponential in the maximum interaction strength and maximum node-degree. We also track down the exact dependence of the rate of tensor recovery on the interaction order $k$, that is allowed to grow with the number of samples and nodes, for both the approaches. Finally, we provide a comparative discussion of the performance of the two approaches based on simulation studies, which also demonstrate the exponential dependence of the tensor recovery rate on the maximum coupling strength.
@@ 1. Introduction
In spite of the presence of a significant amount of literature on matrix structure learning in classical $2$-spin Ising models, to the best of our knowledge, analogous literature for tensor recovery in $k$-spin Ising models is extremely scarce, if at all existent. Only recently, Liu at al. [29] extended the technique and results in [36] to tensor recovery in $k$-spin Ising models. In this paper, we study the interaction screening [38] and the regularized pseudolikelihood methods for tensor recovery in $k$-spin Ising models. As was observed in [5], the regularized pseudolikelihood algorithm performs very poorly for high values of the maximum coupling strength, despite working with a sample size much larger than the logarithm of the number of nodes, and irrespective of the choice of the LASSO tuning parameter. Further, the threshold for the maximum coupling strength, above which this poor performance emerges, is also very near to the critical point for the Ising model on a randomly diluted grid (see [5]). The dependence of the minimum sample size on the maximum coupling strength was not made explicit in [36] or [29], and was incorporated within the proportionality constant. Inspired by [38], in this paper, we work out this dependency for tensor Ising models, and show that a sample size exponential in the maximum coupling strength guarantees consistent tensor recovery. This is a plausible explanation behind the worse behavior of the regularized estimators at low temperatures (high coupling strengths). We also track down the dependence of the minimum sample size and the rate of convergence on the order $k$ of interaction. In the following subsection, we describe the $k$-spin Ising model and the tensor recovery problem in more details.
@@ 1.2. Related Work on Parametric Inference in Ising Models
${\mathbb{P}}(\bm{x})\propto e^{\beta\bm{x}^{\top}\bm{J}\bm{x}+h\sum_{i=1}^{n}x_{i}}\quad(\bm{x}\in\{-1,1\}^{p})$
@@ 1.3. Structure of the Paper
The rest of the paper is organized as follows. In Section 2, we state the main results of this paper. We begin by discussing about a general theory of $\ell^{1}$-penalized $M$-estimators from [1], that will help us prove the main results in this paper. In the same section, we use this general theory to prove theoretical guarantees for the RPLE and the RISE. In Section 3, we provide some simulation studies to demonstrate the comparative performance of these two approaches.
@@ 2. Main Results
In this section, we state and prove the main results of this paper. To begin with, we sketch a general theory about $\ell^{1}$-penalized $M$-estimators from [1], that is crucial in proving our main results,
2.1. A general theory of $\ell^{1}$-penalized $M$-estimators
We use a framework from [1] for general $\ell^{1}$-regularized $M$-estimators to establish that our method works consistently. It turns out that imposing just the following two conditions on the loss function is enough to control the error of the $\ell_{1}$-regularized $M$-estimator:
@@ 2.1. A general theory of $\ell^{1}$-penalized $M$-estimators
The second condition guarantees that the loss function is strongly convex in a conically restricted neighborhood of ($\mathbb{R}^{p-1\choose k-1}$. We will verify Conditions 1 and 2 for the RISE and the RPLE later. With these two conditions in hand, we can finally proceed to bound the error of estimation of the neighborhood parameter $\bm{J}_{r}$. The following proposition (Theorem 1 in [1]) serves as the fundamental result behind our main theorem about the theoretical guarantee of the RPLE:
@@ 2.2. The Regularized Pseudo-Likelihood Estimator
$\|\hat{\bm{J}}_{r,P}-\bm{J}_{r}\|\leqslant\frac{M_{2}\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{{\binom{p-1}{k-1}}}{\varepsilon}}{n}}.$
@@ 2.3. The Regularized Interaction Screening Estimator
Figure 1. Plot of the estimation error (vertical axis) against the sample size $n$ (horizontal axis). The top-left window corresponds to the case $\beta=1$, the top-right window to the case $\beta=1.5$, the bottom-left window to the case $\beta=2$ and the bottom-right window to the case $\beta=2.5$.
@@ 3. Simulation Studies
over a grid of values of $\lambda$, where $\mathcal{L}_{p}$ denotes the objective function, which is $\mathcal{S}$ for the RISE and $\ell$ for the RPLE, and $\hat{J}_{r,\lambda}$ is the estimate corresponding to the penalty parameter $\lambda$.
@@
Figure 2. Plot of the estimation error (vertical axis) against the maximum coupling intensity $\beta$ (horizontal axis) for sample size $n=10^{5}$.
@@
For our numerical experiments, we work with the Ising model on $3-$regular, $3-$uniform hypergraphs on $16$ nodes. In Figure 1, we plot the estimation error of the recovered hypergraphs against the sample size $n$, for four values of the maximum coupling intensity $\beta$, namely $1,1.5,2,2.5$. We observe that the estimation errors are higher for both the RPLE and the RISE for higher values of $\beta$, as expected. Further, the simulations suggest that the RISE performs better than the RPLE in terms of the estimation error, which leaves open the natural question of whether this better performance of the RISE can be demonstrated theoretically. In Figure 2, we plot the estimation error for a fixed sample size $n=10^{5}$ against varying maximum coupling strengths $\beta$. The figure clearly shows an exponentially increasing dependence of the estimation error on $\beta$, as suggested in Theorems 1 and 2.
@@ Appendix A Verification of Conditions 1 and 2 for the Pseudolikelihood
${\mathbb{P}}\left(\left|\frac{\partial\ell(\bm{J}_{r};\mathfrak{X}^{n})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}\right|\geq\frac{{2\sqrt{2}}k!t}{\sqrt{n}}\right)\leq 2e^{-t^{2}}~{}.$
@@
For every $\delta>0$, $\varepsilon>0$ and $n\geq\frac{2}{\delta^{2}}\log\frac{2\binom{p-1}{k-1}^{2}}{\varepsilon}$, we have the following with probability at least $1-\varepsilon$:
@@
thereby completing the proof of Lemma 3. ∎
@@
which completes the proof of Lemma 4. ∎

### v2

@@ Abstract.
In this paper, we study two well known methods of Ising structure learning, namely the pseudolikelihood approach and the interaction screening approach, in the context of tensor recovery in $k$-spin Ising models. We show that both these approaches, with proper regularization, retrieve the underlying hypernetwork structure using a sample size logarithmic in the number of network nodes, and exponential in the maximum interaction strength and maximum node-degree. We also track down the exact dependence of the rate of tensor recovery on the interaction order $k$, that is allowed to grow with the number of samples and nodes, for both the approaches. We then provide a comparative discussion of the performance of the two approaches based on simulation studies, which also demonstrates the exponential dependence of the tensor recovery rate on the maximum coupling strength. Our tensor recovery methods are then applied on gene data taken from the Curated Microarray Database (CuMiDa), where we focussed on understanding the important genes related to hepatocellular carcinoma.
@@ 1. Introduction
In spite of the presence of a significant amount of literature on matrix structure learning in classical $2$-spin Ising models, to the best of our knowledge, analogous literature for tensor recovery in $k$-spin Ising models is extremely scarce, if at all existent. Only recently, Liu at al. [30] extended the technique and results in [37] to tensor recovery in $k$-spin Ising models. In this paper, we study the interaction screening [39] and the regularized pseudolikelihood methods for tensor recovery in $k$-spin Ising models. As was observed in [5], the regularized pseudolikelihood algorithm performs very poorly for high values of the maximum coupling strength, despite working with a sample size much larger than the logarithm of the number of nodes, and irrespective of the choice of the LASSO tuning parameter. Further, the threshold for the maximum coupling strength, above which this poor performance emerges, is also very near to the critical point for the Ising model on a randomly diluted grid (see [5]). The dependence of the minimum sample size on the maximum coupling strength was not made explicit in [37] or [30], and was incorporated within the proportionality constant. Inspired by [39], in this paper, we work out this dependency, and show that a sample size exponential in the maximum coupling strength guarantees consistent tensor recovery. This is a plausible explanation behind the worse behavior of the regularized estimators at low temperatures (high coupling strengths). We also track down the dependence of the minimum sample size and rate of convergence on the order $k$ of interaction. In the following subsection, we describe the $k$-spin Ising model and the tensor recovery problem in more details.
@@ 1.2. Related Work on Parametric Inference in Ising Models
${\mathbb{P}}(\bm{x})\propto e^{\beta\bm{x}^{\top}\bm{J}\bm{x}+h\sum_{i=1}^{n}x_{i}}\quad(x\in\{-1,1\}^{p})$
@@ 1.3. Structure of the Paper
The rest of the paper is organized as follows. In Section 2, we state the main results of this paper. We begin by discussing about a general theory of $\ell^{1}$-penalized $M$-estimators from [1], that will help us prove the main results in this paper. In the same section, we use this general theory to prove theoretical guarantees for the RPLE and the RISE. In Section 3, we provide some simulation studies to demonstrate the comparative performance of these two approaches. In Section 4, we implement these methods on a real-life gene data taken from the Curated Microarray Database (CuMiDa), with the aim of discovering the important genes related to hepatocellular carcinoma. Proofs of some technical results that are not included in the main paper are given in the appendix.
@@ 2. Main Results
In this section, we state and prove the main results of this paper. To begin with, we sketch a general theory about $\ell_{1}$-penalized $M$-estimators from [1], that is crucial in proving our main results,
2.1. A general theory of $\ell_{1}$-penalized $M$-estimators
We use a framework from [1] for general $\ell_{1}$-regularized $M$-estimators to establish that our method works consistently. It turns out that imposing just the following two conditions on the loss function is enough to control the error of the $\ell_{1}$-regularized $M$-estimator:
@@ 2.1. A general theory of $\ell_{1}$-penalized $M$-estimators
The second condition guarantees that the loss function is strongly convex in a conically restricted neighborhood of ($\mathbb{R}^{p-1\choose k-1}$. We will verify Conditions 1 and 2 for the RISE and the RPLE later. With these two conditions and the above assumption in hand, we can finally proceed to bound the error of estimation of the neighborhood parameter $\bm{J}_{r}$. The following proposition serves as the fundamental result begind our main theorem about the theoretical guarantee of the RPLE:
@@ 2.2. The Regularized Pseudo-Likelihood Estimator
$\|\hat{\bm{J}}_{r,P}-\bm{J}_{r}\|_{2}\leqslant\frac{M_{2}\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{{\binom{p-1}{k-1}}}{\varepsilon}}{n}}.$
@@ 3. Simulation Studies
over a grid of values of $\lambda$, where $\mathcal{L}$ denotes the objective function, which is $\mathcal{S}$ for the RISE and $\ell$ for the RPLE, and $\hat{J}_{r,\lambda}$ is the estimate corresponding to the penalty parameter $\lambda$.
(a) $\beta=1$
(b) $\beta=1.5$
(c) $\beta=2$
(d) $\beta=2.5$
@@
Figure 1. Plot of the estimation error (vertical axis) against the sample size $n$ (horizontal axis), for a $3$-uniform, $3$-regular hypergraph on $16$ nodes
@@
For our numerical experiments, we work with the Ising model on $3-$regular, $3-$uniform hypergraphs on $16$ nodes. In Figure 1, we plot the estimation error of the recovered hypergraphs against the sample size $n$, for four values of the coupling intensity $\beta$, namely $1,1.5,2,2.5$. We observe that the estimation errors are higher for both the RPLE and the RISE for higher values of $\beta$, as expected. Further, the simulations suggest that the RISE performs better than the RPLE in terms of the estimation error, which leaves open the natural question of whether this better performance of the RISE can be demonstrated theoretically. In Figure 2, we plot the estimation error for a fixed sample size $n=10^{5}$ against varying coupling strengths $\beta$. The figure clearly shows an exponentially increasing dependence of the estimation error on $\beta$, as suggested in Theorems 1 and 2.
Figure 2. Plot of the estimation error (vertical axis) against the maximum coupling intensity $\beta$ (horizontal axis) for sample size $n=10^{5}$
4. Real Data Analysis
In this section, we implement our methods on gene data from the Curated Microarray Database (CuMiDa)[21], a database composed of 78 handpicked microarray data sets for Homo sapiens that were carefully examined from more than 30,000 microarray experiments from the Gene Expression Omnibus using a rigorous filtering criterion. All data sets were individually submitted to background correction, normalization, sample quality analysis, and were manually edited to eliminate erroneous probes.
The data from the microarray experiment consists of a two-dimensional matrix with genes as rows and samples as columns (coming from different conditions). Each cell in the matrix is a normalized real number indicating how much a gene is expressed in a sample. These expression matrices will usually have thousands of rows and dozens or hundreds of columns. We selected the one particular data set with the largest sample size to facilitate our analysis. The $Liver\_GSE14520\_U133A$ data set contains 357 samples, of which 181 samples have Hepatocellular carcinoma (HCC), and 176 samples are normal. For each sample there are 22279 genes. To reduce the dimension of the data, feature selection was conducted through the Random Forest algorithm. A set of 51 genes was selected as the nodes in the tensor Ising model. The correspondence between the genes and node numbers is given in Appendix D. The values for each node were standardized to binary values $\{-1,+1\}$ by mapping original gene expression values above the mean to $+1$ and values below the mean to $-1$. We then fitted the data of HCC samples and normal samples to a 3-spin Ising model separately, and recovered the corresponding tensors associated to the HCC patients and normal people, using the RISE algorithm.
For both the normal and HCC samples, the fitted hyperedges were then ordered by the absolute value of their coefficients, and the vertices were ordered according to their degrees in the tables above. In Table 1, it shows that the group effects of genes X200910_at, X212551_at and X217165_x_at, and the genes X203554_x_at, X212551_at and X217165_x_at have highest weights in normal samples. From Table 2, we notice that the group effect of genes X210481_s_at, X216025_x_at and X216661_x_at, and genes X201293_x_at, X210481_s_at and X212661_x_at have higher weights in HCC samples. Tables 3 and 4 illustrate that the genes X201453_x_at and X217165_x_at are more important in the samples of normal people, and genes X210481_s_at and X207609_s_at are more related to Hepatocellular carcinoma.
Table 1. Triangles recovered from normal samples ordered by absolute value of edge weights
  (3,33,40) (10,33,40) (13,28,39) (13,28,46) (13,28,38) (6,9,46) (6,9,39) (6,9,38) (40,41,48) 0.3375 0.3375 0.2986 0.2986 0.2986 0.2874 0.2874 0.2874 0.2723
Table 2. Triangles recovered from HCC samples ordered by absolute value of edge weights
  (29,38,39) (5,29,35) (16,29,50) (4,10,23) (14,18,24) (29,36,40) (21,29,42) (14,47,49) (29,31,35) -0.5683 -0.5363 -0.4622 -0.4358 -0.4352 -0.4144 -0.4041 -0.3871 -0.3853
Table 3. Frequency of nodes in hyperedges recovered from normal samples
  Node 6 40 48 33 9 13 28 10 3 Frequency 11 6 4 4 3 3 3 3 3
Table 4. Frequency of nodes in hyperedges recovered from HCC samples
  Node 29 23 14 28 10 47 16 50 Frequency 13 5 4 4 3 3 3 3
@@ Appendix A Verification of Conditions 1 and 2 for the Pseudolikelihood
${\mathbb{P}}\left(\left|\frac{\partial\ell(\bm{J}_{r};\mathfrak{X}^{n})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}\right|\geq\frac{{\color[rgb]{0,0,0}2\sqrt{2}}k!t}{\sqrt{n}}\right)\leq 2e^{-t^{2}}~{}.$
@@
For every $\delta>0$, $\varepsilon>0$ and $n\geq\frac{2}{\delta^{2}}\log\frac{2\binom{p-1}{k-1}^{2}}{\varepsilon}$, we have the following with probability at least $1-\varepsilon$, we have the following:
@@
thereby completing the proof of Lemma 3.
@@
which completes the proof of Lemma 4.
@@ Appendix C Technical Lemmas
Appendix D Correspondence of node numbers and genes
In this section, we provide the correspondence between the node labels and the corresponding gene labels in the following tabular form.
Node Gene
1 X200734_s_at
2 X200882_s_at
3 X200910_at
4 X201128_s_at
5 X201293_x_at
6 X201453_x_at
7 X201656_at
8 X202861_at
9 X203316_s_at
10 X203554_x_at
11 X204428_s_at
12 X204641_at
13 X205019_s_at
14 X205440_s_at
15 X205554_s_at
16 X205866_at
17 X205890_s_at
Node Gene
18 X205911_at
19 X206380_s_at
20 X206453_s_at
21 X206643_at
22 X206680_at
23 X207609_s_at
24 X207804_s_at
25 X207828_s_at
26 X207995_s_at
27 X209219_at
28 X209365_s_at
29 X210481_s_at
30 X211745_x_at
31 X211978_x_at
32 X212149_at
33 X212551_at
34 X212554_at
Node Gene
35 X212661_x_at
36 X213629_x_at
37 X215138_s_at
38 X216025_x_at
39 X216661_x_at
40 X217165_x_at
41 X217319_x_at
42 X217521_at
43 X218002_s_at
44 X218061_at
45 X219787_s_at
46 X220017_x_at
47 X220114_s_at
48 X220148_at
49 X220491_at
50 X220496_at
51 X36829_at
Table 5. Correspondence of node and gene labels.